MITGaard Calendar

beginning of the semester events

Listed below are some of the events happening shorlty that are important for newcomers to the SCA (you don’t have to go to all of them). The stared (*) items are events which do not require any preperation or knowledge of the SCA to participate in. Come to a MITGaard meeting or contact sca@mit.edu to get more information about all SCA events.

* MITGaard meetings are every Wednesday from 6:30 to 7:30 pm in PDR#3^{1} (EX-CEPT for the first one on August 29, which is in 16-168).

* Wednesday, 8/29, 6:30 pm:  Orientation Meeting & Tunic-Making, Part I

An introduction to MITGaard and the SCA. Learn about events, guilds, practices, personas, kingdoms, and how to get more involved. The meeting will be followed by a workshop for making medieval tunics, a good first step in getting into the SCA.

Saturday, 9/1, 11:00 am:  Tunic-Making, Part II (continuation of 8/29)

Sew the tunics begun at the Wednesday tunic-making workshop.

* Wednesday, 9/5, 7:30 pm:  Event Preparation & Dance Practice

You need to come to this meeting to come to the first event of the year, Legends of Chivalry. Afterwards, learn and practice 15th and 16th century dances from Italy, England, and France. No partner or experience is necessary. Practice is every Wednesday, usually in La Sala de Puerto Rico in the Student Center (W20) or in 407 or 491 in the same building (look for signs). Dress is common street clothing.

Weekend of 9/14 - 9/16, travel TBA: Legends of Chivalry

Legends is a camping event at Camp Masopoag set in the year 1217. The event will include armored combat, a hurley (medieval field hockey) competition, archery, weapons throwing, and dancing. You must have garb (medieval clothing) to attend.

Saturday, 9/22, travel TBA:  Falling Leaves

Day-long event at Wellesley College. There will be games, singing, dancing, martial arts, hurley, a play, merchants, and a feast. You must have garb (medieval clothing) to attend.

* Wednesday, 9/26, 6:30:  Chainmail workshop

Learn to make and weave chainmail, a form of armor used since ancient times. Please bring your own needle nose pliers, if you have some; all other materials will be provided. This activity will be held concurrently with MITGaard meetings and will continue throughout the semester.

Saturday, 9/29, travel TBA:  Royal Coronation

Event surrounding the coronation of the East Kingdom’s new King and Queen. Includes Viking long ships, feasting, fighting (tourneys and melees), shopping, archery, and camping. You must have garb (medieval clothing) to attend.

There are many more activities you can get involved in through MITGaard. Come to our meetings to find out more.

1 Private Dining Room #3, on the third floor of the Student Center, W20